OpenStudy (anonymous):

I DARE you find the prime in these set of numbers: 961, 1961, 2961, 3961, 4961, 5961, 6961, 7961, 8961, 9961. I'll give you a medal and follow you if you find the prime. Hint: What's 31^2?

OpenStudy (anonymous):

6961

OpenStudy (anonymous):

Pfft

OpenStudy (anonymous):

6961 is a prime number

OpenStudy (whpalmer4):

@OS31P.vnd I DARE you to find the next prime ending with 961 :-)

ganeshie8 (ganeshie8):

*

OpenStudy (anonymous):

^ agreed

OpenStudy (anonymous):

http://primes.utm.edu/lists/small/10000.txt

OpenStudy (anonymous):

19961

OpenStudy (ikram002p):

ok i dare u to find 10 prime numbers all its digits 1

OpenStudy (ikram002p):

for example 11 is prime all its digits 1 :D

ganeshie8 (ganeshie8):

the dumb way i can think of is to test for divisibility of all primes less than \(\large \sqrt{n}\) I'm not sure how to use the given hint :/

ganeshie8 (ganeshie8):

@KingGeorge

ganeshie8 (ganeshie8):

@oldrin.bataku

OpenStudy (kinggeorge):

I couldn't resist. We know that \(31^2=961\). So 961 is clearly not prime. Furthermore, since the sum of digits is divisible by 3, we know that 2961, 5961, and 8961 aren't prime. Using a similar divisibility test, 9961 is divisible by 7, so that isn't prime either. These are the easy ones. We're left with \[1961,3961,4961,6961,8961\]

OpenStudy (kinggeorge):

A test for 11 tells us that \(4961\) is divisible by 11.

OpenStudy (kinggeorge):

So we only have\(1961,3961,6961\) and \(7961\) left.

ganeshie8 (ganeshie8):

Ahhh looks this problem is designed specifically for this approach ! xD

OpenStudy (kinggeorge):

That method won't work well with 1961. I can't actually find a suitable primality test easy enough to do by hand other than guess and check for the remaining numbers (1961 would take the longest).

OpenStudy (kinggeorge):

I suppose we could use fast exponentiation with 2 (the hardest part will be reducing modulo \(n\)) to see that\[1961,3961,7961\]are all composite since they fail Fermat's little theorem...

OpenStudy (kinggeorge):

For example, write \[1960=2^{10}+2^9+2^8+2^7+2^5+2^3\]Then\[\large\begin{aligned}2^{1960}&\equiv2^{2^{10}+2^9+2^8+2^7+2^5+2^3}\pmod{1961}\\ &\equiv2^{2^{10}}2^{2^9}2^{2^8}2^{2^7}2^{2^5}2^{2^3}\pmod{1961}\\ \end{aligned}\] Then we have that \(2^{2^3}\equiv256\pmod{1961}\), \[2^{2^4}\equiv256^2\equiv823\pmod{1961}\]\[2^{2^5}\equiv823^2\equiv784\pmod{1961}\]\[2^{2^6}\equiv784^2\equiv863\pmod{1961}\]\[2^{2^7}\equiv863^2\equiv-411\pmod{1961}\]\[2^{2^8}\equiv(-411)^2\equiv275\pmod{1961}\]\[2^{2^9}\equiv275^2\equiv1107\pmod{1961}\]\[2^{2^{10}}\equiv1107^2\equiv1785\pmod{1961}\]These are not trivial steps to do by hand, and it probably would be easier to just guess and check. But the we see that\[2^{1960}\equiv-1785*1107*275*411*784*256\equiv1785\not\equiv\pm1\pmod{1961}\]So 1961 doesn't satisfy Fermat's little theorem and must therefore be composite.

OpenStudy (kinggeorge):

It would also have been possible to just exponentiate to 980 and get a similar contradiction. Or perhaps something a bit easier would be the Miller-Rabin test.

OpenStudy (kinggeorge):

You could do similar things for 3961 and 7961, or you could notice that they're divisible by 17 and 19 respectively and therefore not prime. Leaving only 6961.

OpenStudy (kinggeorge):

Or we use the divisibility rules from here. http://en.wikipedia.org/wiki/Divisibility_rule Then 3961 is divisible 17 since 396-5=391 is divisible by 17 since 39-5=34=17*2 is divisible by 17. Also 7961 is divisible by 19 since 796+2=798 is divisible by 19 since 79+16=95 is divisible by 19 since 9+10=19 is divisible by 19. Finally, 1961 is divisible by 37 since 196-11=185=37*5 is divisible by 37.